An air core circular coil has 5000 turns. The inner and outer diameters are 20 and 24 cm

respectively. If a current of 2A is passed through the coil, determine the flux density inside the coil. Show complete solution.


1
Expert's answer
2021-10-07T03:15:12-0400

Find the mean radius:


"r=\frac{D+d}{4}.\\\space\\"

The field is


"B=\mu_0\frac N{2\pi r}I=\frac{2\mu_0NI}{\pi(D+d)},
\\\space\\
B=\frac{2\mu_0·5000·2}{\pi(0.24+0.2)}=0.018\text{ T}."